# Large Language Model-Assisted UAV Operations and Communications: A Multifaceted Survey and Tutorial

### v1

@@ II-C LLM Fine-tuning
Fine-tuning adapts a pretrained LLM to domain-specific tasks using relatively small datasets, significantly reducing data and computational requirements compared to training from scratch. Using representations learned during large-scale pretraining, fine-tuning enables efficient specialization for UAV networks, improving task performance while maintaining generalization. This makes it a practical approach for deploying LLMs under UAV-specific constraints, such as limited computation and dynamic operating environments.
@@
II-D Retrieval-Augmented Generation
RAG offers an effective approach to adapting LLMs to the dynamic and rapidly evolving field of UAV networks by grounding generation in external, up-to-date knowledge. Unlike conventional LLMs that depend solely on static training data, RAG retrieves relevant information at inference time, allowing models to incorporate evolving standards, architectures, and transmission technologies critical to UAV networking. Access to authoritative sources such as 3GPP and IEEE standards, emerging network paradigms, and recent advances in signal processing significantly improve the factual accuracy and contextual relevance of the generated outputs.
By grounding responses in curated domain-specific corpora, RAG reduces hallucinations and outdated information without the prohibitive computational cost of training models from scratch or performing extensive fine-tuning. This makes RAG a practical and scalable approach for leveraging general-purpose LLMs in specialized UAV networking tasks [137]. More importantly, evidence-grounded generation enhances trustworthiness by enabling traceable, policy-aligned, and verifiable decision-making. In safety-critical UAV networks, where accuracy, accountability, and compliance are essential, retrieval-assisted grounding underpins the responsible deployment of AI.
A conventional RAG pipeline, as shown in Fig. 3, typically includes several stages: document indexing in vector databases, query processing and embedding, retrieval and ranking of relevant documents, prompt augmentation with retrieved context, and final response generation. These components together balance retrieval accuracy, response latency, and system scalability. Practical deployment also requires alignment with the application context, secure infrastructure, and governance mechanisms to ensure reliable operation.
@@ II-D Retrieval-Augmented Generation
Despite these successes, conventional RAG suffers from inherent limitations, including fragmented retrieval, redundancy, and limited modeling of inter-document relationships. To address these issues, GraphRAG [91] retrieves structured graph elements rather than isolated text chunks, capturing relational dependencies and improving performance on complex reasoning tasks. Building on this idea, HybridRAG combines vector-based semantic retrieval with graph-based relational retrieval to provide richer grounding for domain-specific applications [100]. Wen et al. [117] further extend HybridRAG by integrating KeywordRAG, VectorRAG, and GraphRAG for multi-UAV mobile edge computing, enabling more reliable formulation of complex carbon-emission optimization problems. Related enterprise-grade RAG architectures developed for robotic systems [133] further illustrate how modular retrieval pipelines can reduce hallucinations by grounding LLM outputs in operational logs, manuals, and safety constraints. However, challenges such as cloud dependence, computational overhead, and sensitivity to data quality remain critical considerations when adapting such RAG-enhanced LLM frameworks to UAV systems.
@@
In summary, RAG improves LLM accuracy by grounding responses in external, up-to-date, domain-specific knowledge. However, it increases integration complexity, depends on high-quality knowledge bases, and incurs variable inference costs based on retrieval strategies and database scale. In contrast, prompt engineering offers a lightweight, low-cost, flexible alternative that requires minimal computational overhead and no additional training. Although it requires expertise and is limited by the model’s context window, prompt engineering enables rapid adaptation to diverse tasks and dynamic wireless network environments.
@@
II-E Prompt Engineering: Strategies for UAV Intelligence
@@
Prompt engineering is the process of designing prompts to guide LLMs in generating accurate and useful outputs. As a reusable and adaptable methodology, prompt engineering supports structured problem-solving across diverse domains, improves understanding of LLM behavior, and enables few-shot learning by optimizing in-context prompts. It also plays an important role in the development of advanced UAV systems. As LLM technologies evolve, prompt engineering is expected to become increasingly important, offering finer control over model outputs and enabling new classes of applications.
Creating effective prompts typically involves several key steps. First, the prompt objective must be clearly defined to ensure that the input structure aligns with the intended task. Second, understanding the model’s capabilities and limitations is essential to ensure compatibility between the prompt and the model’s generation abilities. Third, selecting an appropriate prompt format is crucial, as well-structured and explicit prompts consistently lead to higher-quality responses. Additionally, providing sufficient contextual information helps prevent generic outputs and reduces the risk of misinformation, particularly in less-represented or complex scenarios. Finally, prompts should be iteratively tested and refined according to the model’s responses to improve accuracy and reliability before deployment [84].
Prompting techniques are especially valuable for UAV applications that require real-time decision-making and low-latency responses, such as trajectory optimization and resource allocation. These techniques enable LLMs to interpret sensor data and environmental information directly, generating actionable outputs without extensive retraining. This avoids the significant computational overhead associated with conventional model training. For instance, ICL enables LLMs to perform new UAV-related tasks by providing example demonstrations, with performance strongly influenced by the design and organization of these examples. Similarly, CoT prompting promotes deeper reasoning by guiding the model through intermediate steps, enabling it to handle complex multi-stage UAV tasks such as dynamic obstacle avoidance, swarm coordination, and multi-agent decision-making.
Prompt engineering is particularly well-suited to the constraints of UAV networks for three main reasons. First, it is highly resource-efficient, requiring only a forward pass during inference and eliminating the computational and memory costs associated with backpropagation and gradient updates. This makes it suitable for deployment on resource-constrained network devices. Second, it offers high flexibility: by modifying textual prompts, a single LLM can be quickly adapted to a wide range of network tasks without additional coding or retraining. Finally, prompt engineering enables rapid implementation by leveraging the LLM’s inference capabilities to achieve low response times, a critical feature for low-latency services in dynamic UAV network environments.
The core techniques of prompt engineering, including ICL, CoT prompting, prompt-based planning, and self-refinement, are depicted in Fig. 4 and are discussed below.
II-E1 In-Context Learning - ICL
@@
ICL enables LLMs to adapt to target tasks by conditioning on natural-language instructions and demonstrations provided directly in the prompt, without modifying model parameters. In UASNets, ICL operates by providing contextualized demonstrations, such as prior network states, scheduling decisions, and environmental conditions, that allow the model to exploit pretrained knowledge for real-time adaptation. The effectiveness of ICL depends heavily on demonstration selection and formatting, requiring domain expertise in both UAV networking and LLM behavior.
In practical UAV networking scenarios, such as data-collection scheduling, ICL enables systems to dynamically adapt to time-varying network conditions by incorporating real-time feedback into successive prompts. Sensory information, including queue lengths, channel conditions, and battery levels, can be converted into natural-language task descriptions and combined with historical decisions to guide the LLM. This approach is especially useful in emergency scenarios, such as Search and Rescue (SAR) missions, where conventional DRL-assisted methods require extensive training time and computational resources [30].
Emami et al.[30] propose ICLDC, an LLM-assisted ICL framework for emergency data collection scheduling in UASNets. By deploying an LLM at the network edge to generate scheduling decisions from natural-language descriptions of logged sensory data, ICLDC significantly reduces cumulative packet loss relative to DQN and Maximum Channel Gain baselines, while achieving faster convergence and lower computational overhead. To improve robustness, the framework incorporates a safety verifier to override unsafe decisions and an attack-detection module to mitigate jailbreaking attempts.
Building on this work, Emami et al.[29] introduce FRSICL for real-time flight resource allocation in UAV-assisted wildfire monitoring systems. FRSICL jointly optimizes data-collection schedules and UAV velocities using environmental feedback, including AoI, channel conditions, and UAV locations. Simulation results show that FRSICL outperforms PPO and Nearest-Neighbor baselines, achieving lower average AoI and more stable velocity control. Extending to broader disaster-response scenarios, Emami et al.[28] demonstrate that ICL-assisted LLMs can support real-time adaptive decision-making across multiple UAV functions, including path planning, velocity control, scheduling, and power management, while also revealing potential vulnerabilities to prompt-based attacks.
In multi-UAV post-disaster monitoring scenarios, Emami et al.[27] propose AIC-VDS, an attention-assisted ICL framework for joint scheduling and velocity control. By compressing high-dimensional sensory data (e.g., queue lengths, battery levels, and channel conditions) through attention mechanisms, the framework enables an edge-deployed LLM to prioritize critical sensors and generate task descriptions for real-time optimization. Simulation results indicate that AIC-VDS significantly outperforms Multi-Agent DQN and Maximum Channel Gain baselines, achieving up to 91% reductions in packet loss.
Beyond networking-centric tasks, Ji et al.[51] introduce the LLM-assisted Cognitive Reasoning Framework (LLM-CRF) for human–swarm teaming in disaster SAR missions. By integrating ICL with HITL verification, LLM-CRF interprets high-level multimodal commands, decomposes tasks hierarchically, and coordinates heterogeneous UAV swarms. Simulation results show reductions in task completion time, improvements in mission success rates, and lower operator workload, highlighting the effectiveness of ICL for intuitive and safe human–swarm interaction. Table III provides a summary of ICL Frameworks for Emergency UAV and Sensor Network Applications.
Despite these advantages, ICL-based UAV systems remain vulnerable to jailbreaking attacks that manipulate prompts to bypass safety mechanisms. Such attacks can be broadly categorized into white-box methods, which assume access to internal model information, and black-box methods, which rely solely on crafted prompts and are more realistic in deployment scenarios. Addressing these vulnerabilities is essential for deploying ICL in safety-critical UAV networks.
@@
II-E2 Chain-of-Thought Reasoning
CoT reasoning represents a significant advance in how AI systems handle complex reasoning tasks by explicitly decomposing problems into intermediate steps [116]. Instead of generating answers directly, CoT prompts models to reason step by step, mirroring human problem-solving processes and improving performance on tasks involving mathematics, logic, planning, and multi-step decision-making. By revealing intermediate reasoning, CoT enhances interpretability and explainability, improves generalization, and helps mitigate hallucinations by constraining generation to verifiable logical sequences.
Since its introduction, CoT prompting has evolved from basic paradigms such as Zero-shot CoT and Few-shot CoT to more advanced and scalable frameworks. These include Auto-CoT, which automatically selects effective demonstrations, and Self-Consistency CoT, which aggregates multiple reasoning paths to improve reliability. More recent extensions, such as Tree-of-Thought (ToT) and Graph-of-Thought (GoT), further structure reasoning as trees or graphs, enabling exploration and comparison of alternative solution trajectories. Although these advanced methods introduce additional computational overhead, they are particularly valuable for complex, intent-driven tasks, such as UAV networking, where transparent, logical, and verifiable decision-making is essential [112].
Zhu et al. [138] demonstrate the effectiveness of CoT in UAV swarm decision-making by integrating an LLM with an RL controller. In their framework, CoT prompting enables the LLM to decompose high-level missions into structured sub-tasks and dynamically adjust the RL reward function based on real-time environmental feedback. A RAG module further enhances situational awareness by fusing multimodal sensor data with knowledge retrieved from prior missions. Simulation results in disaster search and formation navigation scenarios show that this approach outperforms conventional multi-agent RL baselines in success rate, cumulative reward, and sample efficiency, highlighting improved adaptability and coordination in dynamic environments.
@@ II-E2 Chain-of-Thought Reasoning
However, ICL and CoT prompting are primarily effective for single-task or narrowly scoped scenarios and may struggle with problems that require extensive information integration or long-horizon, multi-stage reasoning. Many UAV networking tasks involve interdependent subtasks, such as coordinated configuration, planning, and execution across multiple agents. Prompt-based planning addresses these limitations by explicitly generating and organizing structured subtask sequences, enabling LLMs to tackle complex objectives more systematically and at scale.
II-E3 Prompt-Based Planning
@@
This capability is especially important for UAV operations, where tasks such as formation reconfiguration, resource allocation in multi-UAV systems, and coordinated mission planning require long-horizon reasoning across multiple agents in dynamic environments. Therefore, designing effective prompts and planning structures is essential to the successful deployment of LLMs in real-time UAV control and decision-making applications.
Although prompt-based planning enables LLMs to decompose complex objectives into executable subtasks, the initial plans and decisions produced by the model may still be imperfect or suboptimal. Therefore, self-refinement is essential.
@@
Core Objective Learn general language patterns, knowledge, and reasoning abilities from a massive corpus. Adapt a pretrained model’s knowledge to a specific task or domain using a smaller dataset. Guide a pretrained model’s output for a specific task by carefully designing the input prompt. Enhance a pretrained model’s output by retrieving and incorporating relevant, up-to-date information from an external knowledge base.
@@ II-E4 Self-Refinement
Building on this concept, Wang et al. [111] propose a closed-loop framework that significantly enhances the reliability and reasoning capabilities of LLMs for autonomous UAV control. Their approach combines a structured prompt framework, Guidelines, Skill APIs, Constraints, and Examples (GSCE), with an iterative feedback and refinement mechanism. After the LLM generates an initial control program using the GSCE framework, the code is executed within a simulator to produce a UAV trajectory. A dedicated LLM-assisted discriminator then evaluates the trajectory against the task objectives and generates targeted diagnostic feedback if deviations are detected. This feedback is subsequently fed back into the GSCE-guided LLM, prompting code refinement and regeneration. The loop of generation, execution, evaluation, and refinement continues until the task is completed or the predefined iteration limit is reached. Experimental results demonstrate that this closed-loop self-refinement system substantially outperforms one-shot generation approaches, achieving task success rates exceeding 90% by enabling the LLM to learn from execution errors and progressively correct its outputs.
A notable merit of self-refinement is its architectural simplicity, as it relies on a single LLM rather than multiple specialized models. However, this simplicity introduces an important consideration: the effectiveness of self-refinement is strongly dependent on the underlying model’s capabilities. Empirical results indicate that performance gains scale with model size and reasoning ability; for instance, when combined with self-refinement, GPT-4 consistently outperforms GPT-3.5 using the same refinement mechanism.
II-F Key Findings and Insights
The adaptation of LLMs to UAV systems requires carefully balancing performance, computational efficiency, and domain specificity. Pre-training offers broad generalization but remains economically and computationally infeasible for most UAV applications. Fine-tuning provides task specialization but depends on high-quality labeled data and introduces risks of overfitting. In contrast, prompt engineering and RAG emerge as lightweight and flexible approaches that enable task adaptation without parameter updates, making them particularly suitable for dynamic UAV environments.
Prompt engineering shapes model behavior through structured instructions and reasoning strategies such as ICL and CoT, enabling real-time decision support with minimal computational overhead. RAG complements this by grounding generation in external, up-to-date knowledge sources, improving factual accuracy in safety-critical UAV operations. While RAG increases system complexity through knowledge base maintenance and retrieval overhead, it reduces hallucination risks and supports continual adaptation without retraining. Importantly, these adaptation strategies should not be viewed as mutually exclusive. Fine-tuning is advantageous for stable, domain-specific behaviors, whereas RAG and prompt-based methods are better suited for rapidly evolving networking standards and mission contexts. Future LLM-assisted UAV systems will likely integrate these techniques within layered architectures, combining lightweight edge models for real-time control with larger models for strategic reasoning and mission planning.
Despite these advantages, effective deployment remains challenging due to constraints such as limited computational and memory resources on edge devices, latency and bandwidth restrictions, long-context processing requirements, reliability and safety concerns, data scarcity, and the dynamic nature of aerial networks. Addressing these challenges necessitates a combination of complementary deployment strategies, including centralized and on-device inference, collaborative and partitioned model architectures, and advances in memory management, long-context handling, chunking mechanisms, and multimodal reasoning.
@@ II-F Key Findings and Insights
The following provide a comprehensive discussion of these techniques, examining their trade-offs, recent advancements, and future research directions for scaling LLMs to meet the stringent operational demands of UAVs in mission-critical and highly dynamic environments.
II-F1 Edge, On-Device, and Hybrid LLM Deployment
Thanks to their strong generalization and reasoning capabilities, LLMs enable UAVs to interpret complex and dynamic environments, thereby supporting coherent control strategies for navigation, surveillance, and autonomous decision-making. These capabilities make LLMs particularly attractive for mission-critical, real-time applications such as SAR, Intelligence, Surveillance, and Reconnaissance (ISR), infrastructure inspection, and persistent aerial monitoring. However, most existing LLM deployments rely on cloud-based infrastructure, which has significant limitations in terms of latency, bandwidth consumption, and data privacy. Such constraints render cloud-centric approaches impractical, or even infeasible, for delay-sensitive or communication-constrained UAV operations. Integrating LLMs into UAV platforms has the potential to enhance onboard perception and autonomy significantly, but it also introduces substantial computational and communication challenges. Existing deployment strategies can be broadly categorized into four primary paradigms, as summarized in Table V.
Table V: Comparison of LLM inference paradigms for UAV systems—centralized edge, on-device, split, and collaborative device–server—detailing deployment locations, advantages, challenges, and enabling techniques for real-time, resource-constrained aerial applications.
Inference Paradigm Deployment Location Pros Cons Key Techniques
Centralized Edge Inference Edge server / Ground station Supports real-time decision-making; suitable for safety-critical operations Communication-intensive; relies on stable connectivity Token reduction, model compression, quantization, pruning, parameter-efficient fine-tuning
Local Inference (On-Device) UAV hardware Improves privacy; minimizes latency; robust to connectivity issues Lightweight models lack reasoning depth, high computation, and energy consumption, limited memory/storage Deployment of small models (e.g., OPT-125M/350M), model optimization for embedded platforms
Split Inference (Model Partitioning) Initial layers on UAV, deeper layers on edge/cloud Reduces onboard computation; limits data transmission to intermediate representations Requires coordination between UAV and edge/cloud; latency depends on network Partitioning of LLM layers, sending compact intermediate representations instead of raw sensor data
Collaborative Device–Server Inference Lightweight models on UAV, powerful models on server Reduces bandwidth usage; suitable for unreliable communication; leverages local and remote computation Complexity in coordination; preliminary outputs may require verification Speculative decoding, hybrid device–server collaboration, local preliminary output with remote refinement

### v2

@@ II-B LLM Fine-tuning for UAV Applications
Fine-tuning adapts a pretrained LLM to domain-specific tasks using relatively small datasets, significantly reducing data and computational requirements compared to training from scratch. Using representations learned during large-scale pretraining, fine-tuning enables efficient specialization for UAV networks, improving task performance while maintaining generalization. This makes it a practical approach for deploying LLMs under UAV-specific constraints, including limited computational resources and dynamic operating environments.
@@
II-C Retrieval-Augmented Generation for UAVs
RAG offers an effective approach to adapting LLMs to the dynamic and rapidly evolving field of UAV networks by grounding generation in external, up-to-date knowledge. Unlike conventional LLMs that depend solely on static training data, RAG retrieves relevant information at inference time, allowing models to incorporate evolving standards, architectures, and transmission technologies critical to UAV networking. Access to authoritative sources such as 3GPP and IEEE standards, emerging network paradigms, and recent advances in signal processing significantly improve the factual accuracy and contextual relevance of the generated outputs. A conventional RAG pipeline, as shown in Fig. 3, typically includes several stages: document indexing in vector databases, query processing and embedding, retrieval and ranking of relevant documents, prompt augmentation with retrieved context, and final response generation. These components together balance retrieval accuracy, response latency, and system scalability.
@@ II-C Retrieval-Augmented Generation for UAVs
Despite these successes, conventional RAG suffers from inherent limitations, including fragmented retrieval, redundancy, and limited modeling of inter-document relationships. To address these issues, GraphRAG [78] retrieves structured graph elements rather than isolated text chunks, capturing relational dependencies and improving performance on complex reasoning tasks. Building on this idea, HybridRAG combines vector-based semantic retrieval with graph-based relational retrieval to provide richer grounding for domain-specific applications [86]. Wen et al. [99] further extend HybridRAG by integrating KeywordRAG, VectorRAG, and GraphRAG to support multi-UAV mobile edge computing, thereby enabling a more reliable formulation of complex carbon-emission optimization problems.
@@
II-D Prompt Engineering: Strategies for UAV Intelligence
@@
Prompt engineering is the process of designing prompts to guide LLMs in generating accurate and useful outputs. As a reusable and adaptable methodology, prompt engineering supports structured problem-solving across diverse domains, improves understanding of LLM behavior, and enables few-shot learning by optimizing in-context prompts. It also plays an important role in the development of advanced UAV systems. As LLM technologies evolve, prompt engineering is expected to become increasingly important, offering finer control over model outputs and enabling new classes of applications. The core techniques of prompt engineering, including ICL, CoT prompting, prompt-based planning, and self-refinement, are depicted in Fig. 4 and are discussed below.
II-D1 In-Context Learning - ICL
@@
ICL enables LLMs to adapt to target tasks by conditioning on natural-language instructions and demonstrations provided directly in the prompt, without modifying model parameters. In UASNets, ICL operates by providing contextualized demonstrations, such as prior network states, scheduling decisions, and environmental conditions, that allow the model to exploit pretrained knowledge for real-time adaptation. The effectiveness of ICL depends heavily on the selection and formatting of demonstrations, requiring domain expertise in both UAV networking and LLM behavior.
In dynamic UAV networking and emergency scenarios, ICL enables LLMs to incorporate real-time sensory feedback—such as queue lengths, channel conditions, battery levels, AoI, and UAV locations—directly into prompts for adaptive decision-making [29]. A series of works by Emami et al. demonstrate the effectiveness of this paradigm across multiple UAV applications. ICLDC applies ICL to emergency data-collection scheduling in UASNets, reducing packet loss and computational overhead compared with DQN and heuristic baselines [29]. FRSICL extends this idea to joint flight-resource allocation for wildfire monitoring, achieving lower AoI and improved stability compared to PPO-based methods [28]. Further generalization to broader disaster-response tasks—including path planning, scheduling, velocity control, and power management—highlights both performance gains and potential prompt-security vulnerabilities [27]. In multi-UAV coordination, the attention-assisted AIC-VDS framework compresses high-dimensional sensory inputs to support joint scheduling and velocity optimization, significantly outperforming multi-agent DRL baselines [26]. Beyond networking optimization, Ji et al. integrate ICL with human-in-the-loop verification in the LLM-CRF framework for human–swarm teaming, enabling hierarchical task decomposition and coordinated control in SAR missions while improving mission success rates and reducing operator workload [48].
Collectively, these studies indicate that ICL provides a lightweight, adaptive alternative to conventional learning-based optimization, enabling real-time, interpretable, and scalable decision-making across UAV networking, resource management, and swarm coordination tasks.
@@
II-D2 Chain-of-Thought Reasoning
CoT reasoning represents a significant advance in AI systems’ ability to handle complex reasoning tasks by explicitly decomposing problems into intermediate steps [98]. Instead of generating answers directly, CoT prompts models to reason step-by-step, mirroring human problem-solving processes and improving performance on tasks involving mathematics, logic, planning, and multi-step decision-making. By revealing intermediate reasoning, CoT enhances interpretability and explainability, improves generalization, and helps mitigate hallucinations by constraining generation to verifiable logical sequences.
Since its introduction, CoT prompting has evolved from basic paradigms such as Zero-shot CoT and Few-shot CoT to more advanced and scalable frameworks. These include Auto-CoT, which automatically selects effective demonstrations, and Self-Consistency CoT, which aggregates multiple reasoning paths to improve reliability. More recent extensions, such as Tree-of-Thought (ToT) and Graph-of-Thought (GoT), further structure reasoning as trees or graphs, enabling exploration and comparison of alternative solution trajectories. Although these advanced methods introduce additional computational overhead, they are particularly valuable for complex, intent-driven tasks, such as UAV networking, where transparent, logical, and verifiable decision-making is essential [95]. Zhu et al. [118] demonstrate the effectiveness of CoT in UAV swarm decision-making by integrating an LLM with an RL controller. In their framework, CoT prompting enables the LLM to decompose high-level missions into structured sub-tasks and dynamically adjust the RL reward function based on real-time environmental feedback.
@@ II-D2 Chain-of-Thought Reasoning
II-D3 Prompt-Based Planning
@@
This capability is especially important for UAV operations, where tasks such as formation reconfiguration, resource allocation in multi-UAV systems, and coordinated mission planning require long-horizon reasoning across multiple agents in dynamic environments. Therefore, designing effective prompts and planning structures is essential to the successful deployment of LLMs in real-time UAV control and decision-making applications. Although prompt-based planning enables LLMs to decompose complex objectives into executable subtasks, the initial plans and decisions produced by the model may still be imperfect or suboptimal. Therefore, self-refinement is essential.
@@
Core Objective Learn general language patterns, knowledge, and reasoning abilities from a massive corpus. Adapt a pretrained model’s knowledge to a specific task or domain using a smaller dataset. Guide the output of a pretrained model for a specific task by carefully designing the input prompt. Enhance the output of a pretrained model by retrieving and incorporating relevant, up-to-date information from an external knowledge base.
@@ II-D4 Self-Refinement
Building on this concept, Wang et al. [94] propose a closed-loop framework that significantly enhances the reliability and reasoning capabilities of LLMs for autonomous UAV control. Their approach combines a structured prompt framework, Guidelines, Skill APIs, Constraints, and Examples (GSCE), with an iterative feedback and refinement mechanism. After the LLM generates an initial control program using the GSCE framework, the code is executed within a simulator to produce a UAV trajectory. A dedicated LLM-assisted discriminator then evaluates the trajectory against the task objectives and generates targeted diagnostic feedback if deviations are detected. This feedback is subsequently used to guide the GSCE-guided LLM, prompting code refinement and regeneration. The loop of generation, execution, evaluation, and refinement continues until the task is completed or the predefined iteration limit is reached. Experimental results demonstrate that this closed-loop self-refinement system substantially outperforms one-shot generation approaches, achieving task success rates exceeding 90% by enabling the LLM to learn from execution errors and progressively correct its outputs.
II-E Key Findings and Insights
@@
Adapting LLMs to UAV systems requires carefully balancing performance, computational efficiency, and domain specificity. The following provides a comprehensive discussion of these techniques, examining their trade-offs, recent advancements, and future research directions for scaling LLMs to meet the stringent operational demands of UAVs in mission-critical and highly dynamic environments.
II-E1 Edge, On-Device, and Hybrid LLM Deployment
LLMs offer strong reasoning and generalization capabilities that can enhance UAV autonomy in navigation, surveillance, and mission-critical tasks such as SAR and ISR. However, most existing deployments rely on cloud infrastructure, which introduces latency, bandwidth overhead, and privacy concerns, making purely cloud-centric solutions unsuitable for delay-sensitive or communication-constrained UAV operations.